Circular No. 7723

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 R6 (LINEAR-SKIFF) B. A. Skiff, Lowell Observatory, reports his discovery of a comet on

CCD images taken by him with the LONEOS telescope on Sept. 25. The object shows a moderately condensed 15′′ coma and a broad tail ∼ 25′′ toward the west-northwest. T. B. Spahr identified this comet with an object observed on two nights (Sept. 11 and 16, previously linked; m_{2} = 19.4–20.0) and reported as asteroidal in appearance by LINEAR; subsequently LIN-EAR observations from Aug. 19 were also identified. J. G. Ries reports that CCD images obtained with the 0.76-m reflector at McDonald Observatory on Sept. 27.3 UT also show this object to be diffuse.

2001 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.11.41907 1^{h}23^{m}53.^{s}51 −22^{o}48′39′′.6  LINEAR
25.31700 1 17 05.72 −23 21 11.6 17.0 Skiff

All the available astrometry, together with the following orbital elements by B. G. Marsden, appear on MPEC 2001-S55:

T = 2001 Oct. 26.6400 TT  ω = 305.^{o}9837
e = 0.485638  Ω = 70.3559 }2000.0
q = 2.115198 AU  i = 17.3381
a = 4.112277 AU n^{o} = 0.1181900 P = 8.339 years

SUPERNOVA 2001el IN NGC 1448 J. Sollerman, B. Leibundgut, European Southern Observatory; and P.

Lundqvist, Stockholm Observatory, report: “Inspection of a CCD spectrum (range 330–1050 nm) of SN 2001el (cf. IAUC 7720), obtained on Sept. 21 UT with the Very Large Telescope UT2 (+ UVES spectrograph) at Paranal, shows it to be a type-Ia supernova well before maximum. The characteristic Si II absorption at 612 nm appears flat-bottomed. Interstellar absorption lines of Ca H and K, and the lines of Na I D (equivalent widths 0.037, 0.031 nm) are observed at the redshift (1162 km/s) of NGC 1448.”

Additional visual magnitude estimates by L. Monard, Pretoria, S. Africa (cf. IAUC 7721): Sept. 22.091 UT, 13.3; 22.870, 13.2; 23.945, 12.8; 24.121, 12.9; 24.992, 12.7.

2001 September 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green